An airplane propeller is rotating at 1900 rev/min.




(a) Compute the angular velocity in rad/s.




(b) How long in seconds does it take for the propeller to turn through 350 degrees?

1
Expert's answer
2022-03-17T09:35:58-0400

(a) Compute the angular velocity in rad/s.

"\omega=2\pi n\\
=\rm 6.28\: rad*1900*\frac{1}{60\:\rm s}=199\: rad/s"

(b) How long in seconds does it take for the propeller to turn through 350 degrees?

"\Delta t=\frac{\Delta \theta}{\omega}"

"\Delta t=\frac{350^\circ*\frac{\pi\:\rm rad}{180^\circ}}{199\: \rm rad/s}=0.031\:\rm s"